### v1

# StoxLSTM: A Stochastic Extended Long Short-Term Memory Network for Time Series Forecasting

Abstract
The Extended Long Short-Term Memory (xLSTM) network has attracted widespread research interest due to its enhanced capability to model complex temporal dependencies in diverse time series applications. Despite its success, there is still potential to further improve its representational capacity and forecasting performance, particularly on challenging real-world datasets with unknown, intricate, and hierarchical dynamics. In this work, we propose a stochastic xLSTM, termed StoxLSTM, that improves the original architecture into a state space modeling framework by incorporating stochastic latent variables within xLSTM. StoxLSTM models the latent dynamic evolution through specially designed recurrent blocks, enabling it to effectively capture the underlying temporal patterns and dependencies. Extensive experiments on publicly available benchmark datasets from multiple research communities demonstrate that StoxLSTM consistently outperforms state-of-the-art baselines with better robustness and stronger generalization ability.
Keywords Time Series Forecasting  $\cdot$ Generative Model  $\cdot$ xLSTM  $\cdot$ State Space Model  $\cdot$ Variational Inference
1 Introduction
Time series forecasting plays a vital role in numerous real-world applications, including financial market analysis [1, 2], energy systems [3, 4], weather forecasting [5], and healthcare monitoring [6]. Traditional recurrent neural networks (RNNs) [7] and their variants [8, 9], such as Long Short-Term Memory (LSTM) networks [8], have been widely used due to their capacity to model sequential data [10, 11, 12]. However, these models often struggle to capture long-range dependencies and intricate temporal structures present in many real-world time series [13].
The Extended Long Short-Term Memory network (xLSTM) [14] has emerged as a powerful architecture that enhances the conventional LSTM by introducing exponential gating mechanisms and modified memory structures. These improvements enable xLSTM to better capture complex temporal dependencies and maintain stability over long sequences. As a result, xLSTM has demonstrated superior performance in various time series tasks compared to traditional recurrent models [15, 16, 17]. Despite these advances, xLSTM remains fundamentally deterministic, limiting its ability to explicitly represent the latent temporal dynamics and stochasticity inherent in many natural processes. This limitation restricts the model’s expressiveness when dealing with highly nonlinear and uncertain time series data.
Figure 1: The recurrent unit of StoxLSTM. Diamonds denote deterministic variables, while circles denote stochastic variables. (a) shows the stochasticized sLSTM block, while (b) shows the stochasticized mLSTM block. $z_{t}$ represents the latent state obtained by stochastic sampling through the reparameterization process.
To overcome these challenges, we propose a stochastic xLSTM termed StoxLSTM by integrating xLSTM within a specially designed stochastic state space modeling (SSM) framework [18]. Rather than treating RNN and SSM as separate components to be combined [19, 20], our approach integrates latent variable dynamics directly into the recurrent structure, effectively reinterpreting the temporal evolution as a stochastic process parameterized by enhanced xLSTM blocks, as shown in Fig. 1. This formulation allows the model to capture complex hidden temporal patterns and uncertainties that go beyond the capabilities of xLSTM networks and their deterministic variants. By integrating stochastic latent dynamics directly into the recurrent architecture, our approach enhances the representational capacity beyond conventional deterministic frameworks, allowing for more flexible and robust modeling of complex temporal behaviors.
The superiority of the proposed StoxLSTM is verified through extensive experiments across a wide range of publicly available benchmark datasets covering diverse domains such as Electricity, Weather, ETT, Solar, Traffic, PEMS, and ILI [21, 22]. The experimental results demonstrate that StoxLSTM consistently outperforms state-of-the-art baselines across multiple evaluation metrics with better robustness and stronger generalization ability. These results validate the superiority and effectiveness of our stochastic state space reformulation in advancing the modeling power of recurrent architectures for time series forecasting. The main contribution of this work can be concluded as follows:
  • To the best of our knowledge, a stochastic xLSTM, termed StoxLSTM, is proposed for the first time. The StoxLSTM improves the raw deterministic xLSTM architecture by integrating stochastic latent variables into the dynamic recurrent evolutions.
  • Efficient parameterization methods of latent temporal dynamics in StoxLSTM is designed. Through the stochasticized xLSTM blocks, the StoxLSTM is enabled to flexibly capture complex, nonlinear, and stochastic dependencies that inherently exists in many time series analysis applications.
  • Comprehensive experiments on a variety of public avaliable benchmark datasets are conducted. Experimental results verified that StoxLSTM can achieve consistent and significant performance gain over current state-of-the-art methods. The superiority and effectiveness of the stochastic state space reformulation of xLSTM are validated.
2 Related Work
2.1 Time Series Forecasting Models
Time series forecasting aims to predict future values based on historical observations, and has seen significant advances through various modeling paradigms. Among these, Transformer-based models [23] have attracted extensive attention due to their powerful sequence modeling capabilities. Notable variants include Informer [13], Autoformer [24], Crossformer [25], and FEDformer [26], which improve efficiency and accuracy for long-term TSF. However, recent work such as LTSF-Linear [27] challenges the supremacy of Transformers by demonstrating that simple linear models can outperform them on long-term TSF. These linear models [27, 28, 29] better preserve temporal order, effectively input longer look-back lengths without overfitting, and benefit from improved embedding strategies, leading to enhanced interpretability [30]. Inspired by these advances, models like iTransformer [31] and PatchTST [32] incorporate linear transformations along the temporal dimension to enhance Transformer-based forecasting.
In parallel, recurrent architectures remain competitive. xLSTM [14] has recently gained prominence due to its ability to capture complex temporal dependencies. Variants such as xLSTMTime [17], P-sLSTM [15], and xLSTM-Mixer [16] have demonstrated superior TSF performance compared to both Transformer- and linear-based models, highlighting the effectiveness of xLSTM in modeling long-range temporal patterns.
2.2 Generative Modeling for Time Series
Generative modeling approaches have emerged as powerful tools for time series analysis, particularly in capturing uncertainty and complex temporal dynamics beyond deterministic predictions. These methods often introduce latent variables or probabilistic frameworks to model the stochastic nature of real-world time series. Classical statistical models such as ARIMA [33], VAR [34] and SSMs [18] have long been used for probabilistic forecasting by modeling trends, seasonality, and inter-variable dependencies. However, these methods often rely on strong assumptions and are limited in their ability to handle nonlinearities and high-dimensional data commonly found in real-world applications.
Recent advances in deep generative modeling have introduced more flexible frameworks that can better capture complex data distributions without restrictive parametric assumptions. Generative adversarial networks (GANs) [35, 36] and diffusion-based models [37, 38] have demonstrated strong capabilities in modeling intricate temporal patterns and generating realistic synthetic sequences. Among these, deep SSMs [19, 11, 39] stand out by explicitly incorporating latent variables to represent hidden states evolving over time, thereby effectively capturing stochastic temporal dependencies. These deep SSMs typically leverage variational inference techniques [40] to approximate intractable posterior distributions, enabling efficient learning and inference. For instance, ProTran [39] combines SSMs with Transformer architectures to model long-term dependencies in a probabilistic manner, achieving strong performance in uncertainty-aware forecasting tasks.
3 Methodology
3.1 Modeling Time series Forecasting with a Deep SSM
Consider a multivariate time series $x_{1:L+T}^{C}$, where $C$ denotes the number of variables (dimensions), $L$ denotes the look-back length, and $T$ denotes the prediction horizon. Each dimension $x_{1:L+T}^{i}$ can be treated as an individual univariate time series.
For the univariate time series $x_{1:L+T}=\{x_{1},x_{2},\ldots,x_{L},\ldots,x_{L+T}\}$, forecasting with deep learning models involves parameterizing the predictive distribution $p_{\theta}(x_{L+1:L+T}|x_{1:L})$ with network parameters $\theta$. When integrating deep learning with SSMs, latent states $z_{1:L+T}$ are introduced to capture latent dynamics, transforming the predictive distribution into the joint distribution ${p_{\theta}}\left({{x_{1:L+T}},{z_{1:L+T}}|{x_{1:L}}}\right)$.
3.1.1 Dependency Mapping of the Designed SSM
To enable efficient computation and reduce error propagation during forecasting, the generative and inference models in StoxLSTM are designed following a stochastic SSM structure with explicit dependency assumptions, as illustrated in Fig. 2. Specifically, the latent state $z_{t}$ is modeled as a Markov process, capturing the temporal evolution of latent features—meaning $x_{t}$ does not depend on previously generated observations but only on the latent state $z_{t}$ and the known history $x_{1:L}$. Formally, the joint distribution factorizes as:
Figure 2: State space models corresponding to the generative model and the inference models in StoxLSTM.
$\begin{array}[]{l}\quad{p_{\theta}}\left({{x_{1:L+T}},{z_{1:L+T}}|{x_{1:{\rm{L}}}}}\right)\\ =\prod\limits_{t=1}^{L+T}{{p_{\theta}}\left({{x_{t}}|{x_{1:t-1}},{z_{1:t}},{x_{1:L}}}\right)\cdot{p_{\theta}}\left({{z_{t}}|{x_{1:t-1}},{z_{1:t-1}},{x_{1:L}}}\right)}\\ =\prod\limits_{t=1}^{L}{{p_{\theta}}\left({{x_{t}}|{z_{t}},{x_{1:t-1}}}\right)\cdot{p_{\theta}}\left({{z_{t}}|{z_{t-1}},{x_{1:t-1}}}\right)}\cdot\prod\limits_{t=L+1}^{L+T}{{p_{\theta}}\left({{x_{t}}|{z_{t}},{x_{1:L}}}\right)\cdot{p_{\theta}}\left({{z_{t}}|{z_{t-1}},{x_{1:L}}}\right)}\end{array}$ (1)
wherein $\prod\limits_{t=1}^{L}{{p_{\theta}}\left({{x_{t}}|{z_{t}},{x_{1:t-1}}}\right)\cdot{p_{\theta}}\left({{z_{t}}|{z_{t-1}},{x_{1:t-1}}}\right)}$ corresponds to the reconstruction of the known history $x_{1:L}$. While $\prod\limits_{t=L+1}^{L+T}{{p_{\theta}}\left({{x_{t}}|{z_{t}},{x_{1:L}}}\right)\cdot{p_{\theta}}\left({{z_{t}}|{z_{t-1}},{x_{1:L}}}\right)}$ models the evolution of future states and observations based on the known past $x_{1:L}$.
Correspondingly, the approximate posterior of latent states $z_{1:T}$ with network parameters $\varphi$ factorizes as:
${q_{\varphi}}\left({{z_{1:L+T}}|{x_{1:L+T}}}\right)=\prod\limits_{t=1}^{L+T}{{q_{\varphi}}\left({{z_{t}}|{z_{1:t-1}},{x_{1:L+T}}}\right)}=\prod\limits_{t=1}^{L+T}{{q_{\varphi}}\left({{z_{t}}|{z_{t-1}},{x_{1:L+T}}}\right)}$ (2)
3.2 A Stochastic Extend Long Short-term Memory Network
The overall structure of StoxLSTM is illustrated in Fig. 3. StoxLSTM performs channel-independent forecasting for multivariate time series by dividing the multivariate time series into multiple individual univariate sequences. Each univariate sequence $x{(i)}\in\mathbb{R}^{1\times L},i=1,2,\ldots,C$ is processed independently by the StoxLSTM module, and the outputs are concatenated to form the multivariate prediction $\widehat{x}\in\mathbb{R}^{C\times T}$. To address the challenges of modeling long sequences faced by xLSTM [15], each individual sequence is divided into multiple consecutive patches.
Figure 3: Overall framework of StoxLSTM.
3.2.1 Design of the Recurrent Unit in StoxLSTM
By incorporating stochastic latent variables $z_{1:t}$ into the original xLSTM recurrent unit, as shown in Figs. 1 and 2, StoxLSTM follows the SSM that respects temporal causality, Markovianity, and non-autoregressive dynamics. Specifically, at each time step $t$, the latent state $z_{t}$ is modeled as a stochastic variable conditioned on the previous latent state $z_{t-1}$ and the current deterministic hidden state $h_{t}$.
Firstly, StosLSTM (the stochasticized sLSTM block in xLSTM, as shown in Fig. 1(a)) forward pass is:
$\begin{array}[]{ll}{c_{t}}={f_{t}}^{\prime}{c_{t-1}}+{i_{t}}^{\prime}{y_{t}}&\text{cell state}\\ {n_{t}}={f_{t}}^{\prime}{n_{t-1}}+{i_{t}}^{\prime}&\text{normalizer state}\\ {h_{t}}={o_{t}}\widetilde{{h_{t}}}&\text{hidden state}\\ z_{t}\sim\mathcal{N}\big{(}\mu_{\theta}(h_{t},z_{t-1}),\sigma_{\theta}(h_{t},z_{t-1})\big{)}&\text{latent state}\\ {x_{t}}=\varphi\left({w_{z}^{T}{z_{t}}+{r_{z}}{h_{t}}+{b_{z}}}\right)&\text{output gate}\end{array}$
Here, $f_{t}$$^{\prime}$ and $i_{t}$$^{\prime}$ denote the stabilized forget gate and stabilized input gate, respectively. $y_{t}$} represents the cell input, and $o_{t}$ denotes the output gate.
Secondly, StomLSTM (the stochasticized mLSTM block in xLSTM, as shown in Fig. 1(b)) forward pass is:
$\begin{array}[]{ll}{c_{t}}={f_{t}}{c_{t-1}}+{i_{t}}{v_{t}}k_{t}^{T}&\text{cell state}\\ {n_{t}}={f_{t}}{n_{t-1}}+{i_{t}}{k_{t}}&\text{normalizer state}\\ {h_{t}}={o_{t}}\odot{\rm{}}{\widetilde{h}_{t}},\quad{\widetilde{h}_{t}}={C_{t}}{q_{t}}/max\left\{{\left|{n_{t}^{T}{q_{t}}}\right|,1}\right\}&\text{hidden state}\\ z_{t}\sim\mathcal{N}\big{(}\mu_{\theta}(h_{t},z_{t-1}),\sigma_{\theta}(h_{t},z_{t-1})\big{)}&\text{latent state}\\ {x_{t}}=\varphi\left({w_{z}^{T}{z_{t}}+{r_{z}}{h_{t}}+{b_{z}}}\right)&\text{output gate}\end{array}$
Here, $k_{t}$ denotes the key input, $q_{t}$ denotes the query input, and $\odot$ represents element-wise multiplication.
In both StosLSTM and StomLSTM, we assume that the latent state $z_{t}$ follows a Gaussian distribution parameterized by a mean $\mu_{\theta}(h_{t},z_{t-1})$ and variance $\sigma_{\theta}(h_{t},z_{t-1})$, both predicted by neural networks. This probabilistic formulation allows the model to capture the inherent uncertainty and stochasticity present in real-world time series data. By sampling $z_{t}$ from this distribution through the reparameterization block, StoxLSTM is able to represent diverse possible future trajectories, effectively capturing the variability and noise inherent in temporal dynamics. The update of the latent state $z_{t}$ thus follows the conditional distribution $p(z_{t}\mid z_{t-1},x_{1:t-1})$. Similarly, the output $x_{t}$ is generated by applying a linear transformation to $z_{t}$ and $h_{t}$, followed by an activation function $\varphi$. The generation of $x_{t}$ follows the conditional distribution $p(x_{t}\mid z_{t},x_{1:t-1})$.
3.2.2 Generative Model
The generative model evolves according to the factorization in Eq. (1). During preprocessing, zero-padding is applied at both ends of the look-back sequence: padding at the front generates the initial input $x_{0}$ with a length equal to the patch stride $S$, and padding at the end aligns the predicted sequence length $T$. The padded sequence is then segmented into patches $p_{0:N}$, where $N=\lceil\frac{L+T+S-P}{S}\rceil$, with patch size $P$. These patches are embedded as:
${x_{p0:pN}}={\rm{Embedding}}({\rm{patching}}([\underbrace{0,\ldots,0,}_{{S}}{x_{1:L}},\underbrace{0,\ldots,0}_{T}]))$ (3)
The embedded patches $x_{p0:pN}$ fed into the StoxLSTM recurrent unit to generate latent states $z_{1:N+1}$ and time series patches $x_{p1:pN+1}$:
$z_{1:N+1},x_{p1:pN+1}=\mathrm{RecurrentUnit}(x_{p0:pN})$ (4)
Therefore, during the reconstruction phase, the latent states $z_{t}$ and observations $x_{t}$ are generated from the probability distributions ${p_{\theta}}(z_{t}|z_{t-1},x_{1:t-1})$ and ${p_{\theta}}(x_{t}|z_{t},x_{1:t-1})$, respectively. While during the forecasting phase, $z_{t}$ and $x_{t}$ are generated from the distributions ${p_{\theta}}(z_{t}|z_{t-1},x_{1:L})$ and ${p_{\theta}}(x_{t}|z_{t},x_{1:L})$, respectively.
The time series patches $x_{p1:pN+1}$ are then flattened and passed through a linear layer to generate the final prediction $x_{L+1:L+T}$:
$x_{L+1:L+T}=\mathrm{FCL}(\mathrm{Flatten}(x_{p1:pN+1}))$ (5)
where $\mathrm{FCL}(\cdot)$ denotes a fully connected layer.
3.2.3 Inference Model
The inference model follows the factorization in Eq. (2) via a bidirectional recurrent unit [40, 19]. Zero-padding is applied at the front of the sequence $x_{1:L+T}$ with a length equal to the patch stride, followed by patching and embedding:
${x_{p0:pN}}={\rm{Embedding}}({\rm{patching}}([\underbrace{0,\ldots,0,}_{\rm{S}}{x_{1:L+T}}]))$ (6)
The bidirectional recurrent unit processes these embedded patches, producing forward and backward hidden states $h_{1:N+1}$ and $g_{N+1:1}$, which are then used to generate the latent states $z_{1:N+1}$:
${z_{1:N+1}},\_\ =\mathrm{RecurrentUnit_{bidirection}}(x_{p0:pN})$ (7)
The bidirectional recurrent unit enables the latent states $z_{t}$ to be generated from the approximate posterior distribution ${q_{\varphi}}\left({{z_{t}}|{z_{t-1}},{x_{1:L+T}}}\right)$. The role of the inference model is to obtain this approximate posterior during training, while during testing, only the generative model is utilized.
3.2.4 Evidence Lower Bound of StoxLSTM
Following the variational inference framework [19, 39, 41, 20, 42], we derive the Evidence Lower Bound (ELBO) as the training objective for StoxLSTM based on the factorization assumptions in Eqs. (1) and (2):
[…]
Here, $KL(\cdot)$ denotes the KL divergence. This objective encourages the model to reconstruct the observed sequence accurately while maintaining consistency between the inferred latent states and their temporal dynamics. Detailed derivations and implementation details are provided in Appendix A.
3.2.5 Complexity Analysis of StoxLSTM
For a time series $x_{1:L+T}^{C}$, the time complexity of xLSTM is $O(L+T)$, and its space complexity is $O(1)$ [14]. Considering that StoxLSTM adopts patching and channel independence, its time complexity is $O(C(L+T))$ and the space complexity is $O(1)$. Compared to Transformer-based models, StoxLSTM exhibits lower time and space complexity. However, it still has a relatively large number of parameters. During training, both the inference and generative models are required, but during inference, only the generative model is used, reducing computational time and memory overhead.
4 Experiments
4.1 Time Series Forecasting
4.1.1 Datasets
We employ 9 publicly available datasets to evaluate the performance of our proposed StoxLSTM for long-term time series forecasting. These datasets include: Electricity [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014, Weather [^2] https://www.bgc-jena.mpg.de/wetter/, Solar [^3] https://github.com/laiguokun/multivariate-time-series-data, Traffic [^4] http://pems.dot.ca.gov, ILI [^5] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html, and four ETT [^6] https://github.com/zhouhaoyi/ETDataset datasets (ETTm1, ETTm2, ETTh1, ETTh2). Additionally, we utilize 4 public datasets to assess the performance of short-term time series forecasting, which include: PEMS [^7] http://pems.dot.ca.gov (PEMS03, PEMS04, PEMS07, and PEMS08). These datasets have been widely used for benchmarking and are publicly available in [24], [27], [16] and [43]. Detailed dataset information is summarized in Table 1.
Table 1: Dataset detailed descriptions. The dataset size is organized in (Train, Validation, Test).
Datasets Dim Horizon Dataset size Frequency Information Tasks
Weather 21 {96, 192, 336, 720} (36887, 5270, 10539) 10 min Weather Long-term forecasting
Electricity 321 {96, 192, 336, 720} (18412, 2632, 5260) Hourly Electricity Long-term forecasting
Solar-Energy 137 {96, 192, 336, 720} (36792, 5256, 10512) 10 min Electricity Long-term forecasting
ETTh1, ETTh2 7 {96, 192, 336, 720} (8640, 2880, 2880) Hourly Temperature Long-term forecasting
ETTm1, ETTm2 7 {96, 192, 336, 720} (34560, 11520, 11520) 15 min Temperature Long-term forecasting
Traffic 862 {96, 192, 336, 720} (12280, 1756, 3508) 10 min Transportation Long-term forecasting
ILI 170 {24, 36, 48, 60} (676, 97, 193) 7 days Illness Long-term forecasting
PEMS03 358 12 (15724, 5241, 5243) 5 min Transportation Short-term forecasting
PEMS04 307 12 (10195, 3398, 3399) 5 min Transportation Short-term forecasting
PEMS07 883 12 (16934, 5644, 5646) 5 min Transportation Short-term forecasting
PEMS08 170 12 (10713, 3571, 3572) 5 min Transportation Short-term forecasting
4.1.2 StoxLSTM Configuration
StoxLSTM is configured with a consistent set of hyperparameters across all experiments: Patch size $P=56$, Stride $S=24$, model dimension d_model=64d\_model=64, and latent dimension d_latent=16d\_latent=16. Other hyperparameters, including fully connected layer sizes, batch size, learning rate, and dropout rate, are adjusted based on dataset dimensionality and size.
In all experiments, we use mean squared error (MSE) as the reconstruction loss. The loss function of StoxLSTM can be expressed as:
${Loss}\_{{function}}=\text{MSE}\left(x_{1:L+T},\hat{x}_{1:L+T}\right)+\frac{\beta}{d}\sum_{i=1}^{d}\left(\log\frac{\sigma_{p_{i}}}{\sigma_{q_{i}}}-\frac{1}{2}+\frac{\sigma_{p_{i}}^{2}+(\mu_{q_{i}}-\mu_{p_{i}})^{2}}{2\sigma_{q_{i}}^{2}}\right)$
where $x_{1:L+T}$ denotes the ground truth sequence, and $\hat{x}_{1:L+T}$ represents the prediction from StoxLSTM. The second term corresponds to the KL divergence between two Gaussian distributions. Here, $p$ denotes either $p_{\theta}\left(z_{t}|z_{t-1},x_{1:L}\right)$ or $p_{\theta}\left(z_{t}|z_{t-1},x_{1:t-1}\right)$, while $q$ denotes $q_{\varphi}\left(z_{t}|z_{t-1},x_{1:L+T}\right)$. $\mu$ and $\sigma$ represent the mean and standard deviation of the respective distributions. $d$ is the latent dimension d_latentd\_latent, and $i$ indexes the $i$-th dimension of the latent variable. The hyperparameter $\beta$ controls the relative weighting between the reconstruction and KL divergence terms, and is typically set to $\beta=500$ in our experiments. The RAdam optimizer [44] is employed alongside a cosine annealing learning rate scheduler.
4.1.3 Baselines and Experiment Settings.
We select three types of state-of-the-art models as baselines for long-term and short-term time series forecasting:
Transformer-based: Informer [13], Autoformer [24], Crossformer [25], FEDformer [26], PatchTST [32] and iTransformer [31].
Linear-based: Dlinear [27] and TimeMixer++ [28, 29].
xLSTM-based: xLSTMTime [17], P-sLSTM [15] and xLSTM-Mixer [16].
All models are evaluated under consistent experimental conditions.
For long-term forecasting, the prediction horizons for the ILI dataset are set to $T\in\{24,36,48,60\}$, with StoxLSTM’s look-back window fixed at $L=96$. For all other datasets, the prediction horizons are $T\in\{96,192,336,720\}$, and the look-back window of StoxLSTM is set to $L=336$. For short-term forecasting, all models are set to $T=12$ and $L=96$. Baseline results and look-back length $L$ settings of long-term forecasting are primarily sourced from PatchTST [32], xLSTM-Mixer [16] and TimeMixer++ [29], supplemented by our own experiments where results are unavailable. Evaluation metrics for long-term forecasting include Mean Squared Error (MSE) and Mean Absolute Error (MAE), while short-term forecasting is assessed using MAE, Mean Absolute Percentage Error (MAPE), and Root Mean Squared Error (RMSE). Experiments are conducted on an NVIDIA RTX 3090 GPU (24GB) and 3 NVIDIA A800 GPUs (80GB).
4.1.4 Metrics
To ensure maximum comparability with prior work [32, 16, 29], we adopt the commonly used evaluation metrics for all models. Specifically, long-term forecasting is assessed using mean absolute error (MAE) and MSE. Short-term forecasting is evaluated with MAE, root mean squared error (RMSE), and mean absolute percentage error (MAPE). These metrics are averaged across all variables and calculated as follows:
${\rm{MAE}}\left({\textbf{y},\widehat{\textbf{y}}}\right)=\frac{1}{{{T}}}\sum\limits_{i=1}^{T}{\left|{{y_{i}}-{{\hat{y}}_{i}}}\right|}\qquad{\rm{MAPE}}\left({\textbf{y},\widehat{\textbf{y}}}\right)=\frac{{100}}{{{T}}}\sum\limits_{i=1}^{T}{\frac{{\left|{{y_{i}}-{{\hat{y}}_{i}}}\right|}}{{\left|{{y_{i}}}\right|+\varepsilon}}}$
${\rm{MSE}}\left({\textbf{y},\widehat{\textbf{y}}}\right)=\frac{1}{{{T}}}\sum\limits_{i=1}^{T}{{{\left({{y_{i}}-{{\widehat{y}}_{i}}}\right)}^{2}}}\qquad{\rm{RMSE}}\left({\textbf{y},\widehat{\textbf{y}}}\right)=\sqrt{\frac{1}{{{T}}}\sum\limits_{i=1}^{T}{{{\left({{y_{i}}-{{\widehat{y}}_{i}}}\right)}^{2}}}}$
wherein y denotes the ground truth, $\widehat{\textbf{y}}$ denotes the predictions, and $\varepsilon$ is a small constant for numerical stability. Additionally, the Continuous Ranked Probability Score (CRPS) evaluation metric is detailed in Appendix B.1.
4.1.5 Results
Table 2 and Table 3 present the results of long- and short-term time series forecasting, respectively. In general, StoxLSTM consistently outperforms all baseline models in nearly all datasets and prediction horizons, achieving the lowest MSE, MAE, MAPE, and RMSE.
Table 2: Long-term time series forecasting results. The best results are in bold and the second best are in underline.
Models StoxLSTM xLSTM-Mixer xLSTMTime P-sLSTM iTransformer PatchTST FEDformer Informer Autoformer Dlinear TimeMixer++
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather
96 0.113 0.168 0.143 0.184 0.144 0.187 0.149 0.208 0.174 0.214 0.149 0.198 0.217 0.296 0.300 0.384 0.266 0.336 0.176 0.237 0.155 0.205
192 0.141 0.207 0.186 0.226 0.192 0.236 0.197 0.256 0.221 0.254 0.194 0.241 0.276 0.336 0.598 0.544 0.307 0.367 0.220 0.282 0.201 0.245
336 0.162 0.225 0.236 0.266 0.237 0.272 0.249 0.297 0.278 0.296 0.245 0.282 0.339 0.380 0.578 0.523 0.359 0.395 0.265 0.319 0.237 0.265
720 0.196 0.255 0.310 0.323 0.313 0.326 0.320 0.350 0.358 0.347 0.314 0.334 0.403 0.428 1.059 0.741 0.419 0.428 0.323 0.362 0.312 0.334
Electricity
96 0.117 0.223 0.126 0.218 0.128 0.221 0.130 0.226 0.148 0.240 0.129 0.222 0.193 0.308 0.274 0.368 0.201 0.317 0.140 0.237 0.135 0.222
192 0.136 0.242 0.144 0.235 0.150 0.243 0.148 0.243 0.162 0.253 0.147 0.240 0.201 0.315 0.296 0.386 0.222 0.334 0.153 0.249 0.147 0.235
336 0.144 0.253 0.157 0.250 0.166 0.259 0.165 0.262 0.178 0.269 0.163 0.259 0.214 0.329 0.300 0.394 0.231 0.338 0.169 0.267 0.164 0.245
720 0.159 0.270 0.183 0.276 0.185 0.276 0.199 0.293 0.225 0.317 0.197 0.290 0.246 0.355 0.373 0.439 0.254 0.331 0.203 0.301 0.212 0.310
Solar
96 0.098 0.179 0.227 0.313 0.241 0.306 0.167 0.232 0.203 0.237 0.167 0.224 0.287 0.383 0.200 0.247 0.456 0.446 0.289 0.377 0.171 0.231
192 0.124 0.212 0.245 0.328 0.265 0.331 0.180 0.241 0.233 0.261 0.189 0.251 0.278 0.364 0.220 0.251 0.588 0.561 0.319 0.397 0.218 0.263
336 0.139 0.224 0.255 0.330 0.267 0.336 0.190 0.248 0.248 0.273 0.193 0.252 0.319 0.397 0.260 0.287 0.595 0.588 0.352 0.415 0.212 0.269
720 0.169 0.241 0.257 0.324 0.261 0.324 0.196 0.249 0.249 0.275 0.204 0.269 0.319 0.402 0.244 0.301 0.733 0.633 0.356 0.412 0.212 0.270
ETTh1
96 0.342 0.398 0.359 0.386 0.368 0.395 0.381 0.405 0.386 0.405 0.370 0.400 0.376 0.419 0.865 0.713 0.449 0.459 0.375 0.399 0.361 0.403
192 0.370 0.411 0.402 0.417 0.401 0.416 0.420 0.431 0.441 0.436 0.413 0.429 0.420 0.448 1.008 0.792 0.500 0.482 0.405 0.416 0.416 0.441
336 0.379 0.419 0.408 0.429 0.422 0.437 0.456 0.458 0.487 0.458 0.422 0.440 0.459 0.465 1.107 0.809 0.521 0.496 0.439 0.443 0.430 0.434
720 0.381 0.422 0.419 0.448 0.441 0.465 0.516 0.512 0.503 0.491 0.447 0.468 0.506 0.507 1.181 0.865 0.514 0.512 0.472 0.490 0.467 0.451
ETTh2
96 0.238 0.307 0.267 0.329 0.273 0.333 0.320 0.378 0.297 0.349 0.274 0.337 0.346 0.388 3.755 1.525 0.358 0.397 0.289 0.353 0.276 0.328
192 0.268 0.340 0.338 0.375 0.340 0.378 0.422 0.442 0.380 0.400 0.341 0.382 0.429 0.441 5.602 1.931 0.456 0.452 0.383 0.418 0.342 0.379
336 0.284 0.356 0.367 0.401 0.373 0.403 0.523 0.494 0.428 0.432 0.329 0.384 0.496 0.487 4.721 1.835 0.482 0.486 0.448 0.465 0.346 0.398
720 0.316 0.384 0.388 0.424 0.398 0.430 0.946 0.669 0.427 0.445 0.379 0.422 0.463 0.474 3.647 1.625 0.515 0.511 0.605 0.551 0.392 0.415
ETTm1
96 0.223 0.310 0.275 0.328 0.286 0.335 0.292 0.343 0.334 0.368 0.293 0.346 0.379 0.419 0.672 0.571 0.505 0.475 0.299 0.343 0.310 0.340
192 0.261 0.339 0.319 0.354 0.329 0.361 0.329 0.369 0.377 0.391 0.333 0.370 0.426 0.441 0.795 0.669 0.553 0.496 0.335 0.365 0.348 0.362
336 0.288 0.357 0.353 0.374 0.358 0.379 0.362 0.391 0.426 0.420 0.369 0.392 0.445 0.459 1.212 0.871 0.621 0.537 0.369 0.386 0.376 0.391
720 0.327 0.384 0.409 0.407 0.416 0.411 0.421 0.424 0.491 0.459 0.416 0.420 0.543 0.490 1.166 0.823 0.671 0.561 0.425 0.421 0.440 0.423
ETTm2
96 0.125 0.224 0.157 0.244 0.164 0.250 0.238 0.302 0.180 0.264 0.166 0.256 0.203 0.287 0.365 0.453 0.255 0.339 0.167 0.260 0.170 0.245
192 0.164 0.259 0.213 0.285 0.218 0288 0.337 0.376 0.250 0.309 0.223 0.296 0.269 0.328 0.533 0.563 0.281 0.340 0.224 0.303 0.229 0.291
336 0.192 0.285 0.269 0.322 0.271 0.322 0.444 0.438 0.311 0.348 0.274 0329 0.325 0.366 1.363 0.887 0.339 0.372 0.281 0.342 0.303 0.343
720 0.277 0.349 0.351 0.377 0.361 0.380 0.488 0.482 0.412 0.407 0.362 0.385 0.421 0.415 3.379 1.388 0.422 0.419 0.397 0.421 0.373 0.399
Traffic
96 0.333 0.258 0.351 0.236 0.358 0.242 0.576 0.332 0.395 0.268 0.360 0.249 0.587 0.366 0.719 0.391 0.613 0.388 0.410 0.282 0.392 0.253
192 0.426 0.311 0.377 0.241 0.378 0.253 0.595 0.342 0.417 0.276 0.379 0.256 0.604 0.373 0.696 0.379 0.616 0.382 0.423 0.287 0.402 0.258
336 0.449 0.320 0.394 0.250 0.392 0.261 0.604 0.345 0.433 0.283 0.392 0.264 0.621 0.383 0.777 0.420 0.622 0.337 0.436 0.296 0.428 0.263
720 0.469 0.336 0.439 0.283 0.434 0.287 0.630 0.357 0.467 0.302 0.432 0.286 0.626 0.382 0.864 0.472 0.660 0.408 0.466 0.315 0.441 0.282
ILI
24 0.930 0.583 1.351 0.707 1.514 0.694 3.964 1.309 1.834 0.883 1.319 0.754 3.228 1.260 4.657 1.449 2.906 1.182 2.215 1.081 1.811 0.823
36 1.181 0.669 1.408 0.712 1.519 0.722 4.095 1.330 1.742 0.870 1.579 0.870 2.679 1.080 4.650 1.463 2.585 1.038 1.963 0.963 1.763 0.835
48 1.283 0.738 1.434 0.721 1.500 0.725 4.097 1.326 1.996 0.815 1.553 0.815 2.622 1.078 5.004 1.542 3.024 1.145 2.130 1.024 1.705 0.818
60 1.385 0.769 1.512 0.737 1.418 0.715 4.355 1.384 1.806 0.788 1.470 0.788 2.857 1.157 5.071 1.543 2.761 1.114 2.368 1.096 1.708 0.839
Wins 33 27 1 6 1 1 0 0 0 0 2 0 0 0 0 0 0 0 0 0 0 3
Table 3: Short-term time series forecasting results. The best results are in bold and the second best are in underline. The prediction length is set to 12.
Models StoxLSTM xLSTM-Mixer P-sLSTM iTransformer PatchTST Crossformer Autoformer FEDformer Dlinear TimeMixer TimeMixer++
PEMS03
MAE 9.74 15.71 17.15 16.72 18.95 15.64 18.08 19.00 19.70 14.63 13.99
MAPE 9.59 14.92 16.56 15.81 17.29 15.74 18.75 18.75 18.35 11.54 13.43
RMSE 14.65 24.82 27.68 27.81 30.15 25.56 27.82 30.05 32.35 23.28 24.03
PEMS04
MAE 14.59 24.61 23.01 21.81 24.86 20.38 25.00 26.51 24.62 19.21 17.46
MAPE 9.25 15.48 14.44 13.42 16.65 12.84 16.70 16.76 16.12 12.53 11.34
RMSE 23.48 38.92 35.92 33.91 40.46 32.41 38.02 41.82 39.51 30.92 28.73
PEMS07
MAE 14.36 27.98 23.54 23.01 27.87 22.79 26.92 27.92 28.65 20.57 18.38
MAPE 6.32 12.20 9.78 10.02 12.69 9.41 11.83 12.29 12.15 8.62 7.32
RMSE 22.34 43.92 37.70 35.56 42.56 35.61 40.60 42.29 45.02 33.59 31.75
PEMS08
MAE 12.55 16.56 17.20 17.94 20.35 17.38 20.47 20.56 20.26 15.22 13.81
MAPE 7.79 10.24 10.87 10.93 13.15 10.80 12.27 12.41 12.09 9.67 8.21
RMSE 19.45 26.56 27.32 27.88 31.04 27.34 31.52 32.97 32.38 24.26 23.62
In long-term forecasting tasks, xLSTM-based models generally demonstrate superior performance compared to Transformer-based and Linear-based models, highlighting the strength of recurrent architectures like xLSTM in capturing long-range temporal dependencies. Among these, StoxLSTM achieves the best results. For instance, on the Weather dataset, StoxLSTM attains an average MSE of 0.161 across different prediction horizons, which is 26.5% lower than the xLSTM-Mixer’s average MSE of 0.219. Similarly, on the Electricity dataset, StoxLSTM achieves an average MSE of 0.139, outperforming xLSTM-Mixer’s 0.153 by 9.1%, further demonstrating its consistent advantage. Notably, StoxLSTM maintains this leading performance across other datasets such as Solar, ETTh1, and ETTm1, reflecting its robustness and generalizability in diverse long-term forecasting scenarios.
In short-term forecasting, although Linear-based models show competitive results, StoxLSTM still surpasses them by a significant margin. Specifically, across four traffic datasets (PEMS03, PEMS04, PEMS07, and PEMS08), StoxLSTM achieves an average MAE of 12.81, outperforming the best Linear-based model, TimeMixer++, which records an MAE of 15.91, representing a 19.49% improvement. Furthermore, StoxLSTM attains an average MAPE of 8.24, which is 18.25% lower than TimeMixer++’s 10.08. In terms of RMSE, StoxLSTM achieves 19.98 on average, outperforming TimeMixer++’s 27.03 by 26.08%. These consistent improvements across multiple metrics and datasets underscore StoxLSTM’s effectiveness in capturing short-term temporal patterns with high accuracy. Some of the forecasting results are shown in Figs. 4 to 6.
Figure 4: The real and forecasting sequences of the last 6 out of 321 time series from the electricity dataset are shown. The orange line denotes the forecasted values generated by StoxLSTM, while the blue line represents the ground truth. The prediction horizon is set to 336 time steps.
Figure 5: The real and forecasting sequences of the last 6 out of 862 time series from the Traffic dataset are shown. The orange line denotes the forecasted values generated by StoxLSTM, while the blue line represents the ground truth. The prediction horizon is set to 96 time steps.
Figure 6: The real and forecasting sequences of the last 6 out of 170 time series from the electricity dataset are shown. The orange line denotes the forecasted values generated by StoxLSTM, while the blue line represents the ground truth. The prediction horizon is set to 48 time steps.
Overall, these results demonstrate that StoxLSTM consistently delivers state-of-the-art accuracy across both short- and long-term forecasting tasks, validating its effectiveness and robustness in diverse time series forecasting scenarios. Supplementary experimental details are provided in B.2.
4.2 Model Analysis
4.2.1 Ablation Study.
We study the effect of patching combined with channel independence (P + CI), series decomposition (SD), stochastic latent variables (stochastic), and the SSM structure on forecasting performance, with results summarized in Table 4. P-sLSTM is employed as a baseline to compare against the SSM design integrated within StoxLSTM. For the model without patching and channel independence, the look-back length is reduced to 96 steps (compared to 336 steps for the other models) to prevent overfitting and reduce training time and memory consumption. Additionally, xLSTM-Mixer is included as the state-of-the-art benchmark for xLSTM-based models.
Table 4: Ablation studies on StoxLSTM are conducted with 5 variants: (a) original StoxLSTM; (b) StoxLSTM without stochastic latent variables, where the latent state $z_{t}$ is generated deterministically without random Gaussian sampling (w/o stochastic); (c) StoxLSTM without patching and channel independence methods (w/o P + CI); (d) StoxLSTM without the series decomposition block (w/o SD); and (e) P-sLSTM, which incorporates P + CI and SD but does not apply the SSM enhancement to the xLSTM architecture.
Models StoxLSTM w/o stochastic w/o P + CI w/o SD P-sLSTM xLSTM-Mixer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Weather
96 0.113 0.168 0.113 0.169 0.106 0.163 0.121 0.180 0.149 0.208 0.143 0.184
192 0.141 0.207 0.139 0.202 0.128 0.188 0.144 0.209 0.197 0.256 0.221 0.254
336 0.162 0.225 0.163 0.226 0.156 0.220 0.166 0.230 0.249 0.297 0.236 0.266
720 0.196 0.255 0.198 0.260 0.206 0.267 0.212 0.266 0.320 0.350 0.310 0.323
ETTm1
96 0.223 0.310 0.229 0.315 0.304 0.375 0.253 0.336 0.292 0.343 0.275 0.184
192 0.261 0.339 0.265 0.340 0.494 0.461 0.275 0.351 0.329 0.369 0.319 0.354
336 0.288 0.357 0.291 0.358 0.592 0.499 0.290 0.360 0.362 0.391 0.353 0.370
720 0.327 0.384 0.329 0.385 0.645 0.526 0.328 0.385 0.421 0.424 0.409 0.407
Electricity
96 0.117 0.223 0.120 0.229 0.357 0.421 0.151 0.260 0.130 0.226 0.126 0.218
192 0.136 0.242 0.136 0.244 0.454 0.498 0.152 0.261 0.148 0.243 0.144 0.235
336 0.144 0.253 0.146 0.255 0.621 0.608 0.158 0.268 0.165 0.262 0.157 0.250
720 0.159 0.270 0.161 0.274 0.699 0.668 0.162 0.278 0.199 0.293 0.183 0.276
Experimental results clearly demonstrate that all components—series decomposition, patching, channel independence, stochastic latent variables, and the SSM structure—contribute positively to the overall performance of StoxLSTM. Focusing first on the SSM modeling aspect, both the original StoxLSTM and its variants without stochastic latent variables (w/o Stochastic) and without series decomposition (w/o SD) consistently outperform P-sLSTM across all experiments, achieving lower MSE and MAE. This confirms that the state space model enhancement to the xLSTM architecture plays a crucial role in improving forecasting performance. Moreover, compared to the w/o Stochastic variant, StoxLSTM with stochastic latent variables generally achieves better and more stable prediction results, indicating that modeling latent states as random variables better captures the underlying temporal dependencies in time series data and enhances robustness. In contrast, the model variant without patching and channel independence exhibits notable instability: while it achieves excellent results in certain scenarios, its performance deteriorates significantly in others. This suggests that patching and channel independence not only improve robustness but also enable the model to effectively handle longer look-back lengths under constrained computational and memory resources, thereby reducing the risk of overfitting.
4.2.2 Representational Capacity of the Latent States.
[…]
Figure 7: Heatmaps of four tensor groups—$x$, $x_{p}$, $z$, and hh—from StoxLSTM when predicting a specific instance in the Electricity dataset. Different colors indicate different values, with brighter colors corresponding to larger values.
Within the inference model, values of $x_{p1:pN}$ and $g_{1:N}$ exhibit limited variation across patches but display noticeable differences across dimensions within each patch. The latent states effectively capture this pattern: both $z\_mean_{1:N}$ and $z\_logvar_{1:N}$ remain relatively stable across patches while preserving significant variability across dimensions. This suggests that the latent representation encodes consistent temporal patterns while retaining important feature-level distinctions.
Notably, through variational inference, the latent states learned by the generative model similarly reflect these characteristics—showing near-uniformity across patches alongside meaningful variation across dimensions within each patch. Such a well-structured latent representation, combined with the hidden states $h_{1:N}$, offers a comprehensive and informative summary of the time series, thereby supporting accurate and robust forecasting.
As demonstrated in Fig. 7, the variational inference training converges effectively, and the latent states successfully capture the underlying temporal features of the data. This visualization not only validates the model’s ability to learn meaningful representations but also highlights the interpretability of the latent space in modeling complex temporal dynamics.
5 Conclusion
In this paper, we propose StoxLSTM, a novel stochastic variant of the original xLSTM architecture. By integrating a designed SSM, StoxLSTM effectively captures the dynamic evolution of time series data. Our model demonstrates a strong capability in both long- and short-term forecasting tasks. Extensive experiments on multiple publicly available benchmark datasets show that StoxLSTM consistently outperforms existing state-of-the-art methods in terms of prediction accuracy and robustness. These results validate the effectiveness of combining the xLSTM framework with the SSM design to enhance temporal modeling. Future work will focus on further optimizing the model architecture and improving computational efficiency.
6 Acknowledgement
This work was supported by the National Natural Science Foundation (NSFC) of China under Grant 62301031.
Appendix A ELBO Derivation of StoxLSTM
The KL divergence between the approximate posterior probability ${q_{\varphi}}\left({{z_{1:L+T}}|{x_{1:L+T}}}\right)$} and the posterior probability $p\left({{z_{1:L+T}}|{x_{1:L+T}}}\right)$} in StoxLSTM is derived as follows:
${\rm{KL}}\left({q_{\varphi}\left(z_{1:L+T}|x_{1:L+T}\right)\|p\left(z_{1:L+T}|x_{1:L+T}\right)}\right)$
$=\,$ $\mathbb{E}_{q_{\varphi}\left(z_{1:L+T}|x_{1:L+T}\right)}\left[\log q_{\varphi}\left(z_{1:L+T}|x_{1:L+T}\right)-\log p\left(z_{1:L+T}|x_{1:L+T}\right)\right]$
$=\,$ $\mathbb{E}_{q_{\varphi}\left(z_{1:L+T}|x_{1:L+T}\right)}\left[\log q_{\varphi}\left(z_{1:L+T}|x_{1:L+T}\right)-\log p\left(z_{1:L+T}|x_{1:L+T},x_{1:L}\right)\right]$
$=\,$ $\mathbb{E}_{q_{\varphi}\left(z_{1:L+T}|x_{1:L+T}\right)}\left[\log q_{\varphi}\left(z_{1:L+T}|x_{1:L+T}\right)-\log p\left(x_{1:L+T},z_{1:L+T}|x_{1:L}\right)+\log p\left(x_{1:L+T}|x_{1:L}\right)\right]$
$=\,$ $\mathbb{E}_{q_{\varphi}\left(z_{1:L+T}|x_{1:L+T}\right)}\left[\log q_{\varphi}\left(z_{1:L+T}|x_{1:L+T}\right)-\log p\left(x_{1:L+T},z_{1:L+T}|x_{1:L}\right)\right]+\log p\left(x_{1:L+T}|x_{1:L}\right)$
Then, ${q_{\varphi}}\left({{z_{1:L+T}}|{x_{1:L+T}}}\right)$} and $p\left({{x_{1:L+T}},{z_{1:L+T}}|{x_{1:L}}}\right)$} can be factorized over time according to Eq. (1) and (2):
$\mathbb{E}_{q_{\varphi}(z_{1:L+T}|x_{1:L+T})}\left[\log q_{\varphi}(z_{1:L+T}|x_{1:L+T})-\log p(x_{1:L+T},z_{1:L+T}|x_{1:L})\right]+\log p(x_{1:L+T}|x_{1:L})$
$=\,$ $\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}(z_{1:L+T}|x_{1:L+T})}\left[\log q_{\varphi}(z_{t}|z_{1:t-1},x_{1:L+T})-\log p_{\theta}(z_{t}|x_{1:t-1},z_{1:t-1},x_{1:L})-\log p_{\theta}(x_{t}|x_{1:t-1},z_{1:t},x_{1:L})\right]$
$\quad+\log p(x_{1:L+T}|x_{1:L})$
$=\,$ $\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}(z_{1:L+T}|x_{1:L+T})}\left[\log q_{\varphi}(z_{t}|z_{1:t-1},x_{1:L+T})-\log p_{\theta}(z_{t}|x_{1:t-1},z_{1:t-1},x_{1:L})\right]$
$\quad-\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}(z_{1:L+T}|x_{1:L+T})}\left[\log p_{\theta}(x_{t}|x_{1:t-1},z_{1:t},x_{1:L})\right]+\log p(x_{1:L+T}|x_{1:L})$
$=\,$ $\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}(z_{t+1:L+T}|x_{1:L+T})}\Bigg{[}\mathbb{E}_{q_{\varphi}(z_{1:t-1}|x_{1:L+T})}\Bigg{[}\mathbb{E}_{q_{\varphi}(z_{t}|z_{1:t-1},x_{1:L+T})}\Big{[}\log q_{\varphi}(z_{t}|z_{1:t-1},x_{1:L+T})-\log p_{\theta}(z_{t}|x_{1:t-1},z_{1:t-1},x_{1:L})\Big{]}\Bigg{]}\Bigg{]}$
$\quad-\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}(z_{t+1:L+T}|x_{1:L+T})}\Bigg{[}\mathbb{E}_{q_{\varphi}(z_{1:t-1}|x_{1:L+T})}\Bigg{[}\mathbb{E}_{q_{\varphi}(z_{t}|z_{1:t-1},x_{1:L+T})}\Big{[}\log p_{\theta}(x_{t}|x_{1:t-1},z_{1:t},x_{1:L})\Big{]}\Bigg{]}\Bigg{]}$
$\quad+\log p(x_{1:L+T}|x_{1:L})$
Moreover, based on the following expression for the KL divergence:
$\mathrm{KL}$ $\left(q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)\,\|\,p_{\theta}\left(z_{t}|x_{1:t-1},z_{1:t-1},x_{1:L}\right)\right)$
$=\,$ $\mathbb{E}_{q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)}\Big{[}\log q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)-\log p_{\theta}\left(z_{t}|x_{1:t-1},z_{1:t-1},x_{1:L}\right)\Big{]}$
The above expression can be simplified as follows:
$\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}\left(z_{t+1:L+T}|x_{1:L+T}\right)}\Bigg{[}\mathbb{E}_{q_{\varphi}\left(z_{1:t-1}|x_{1:L+T}\right)}\Bigg{[}\mathbb{E}_{q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)}\Big{[}\log q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)-\log p_{\theta}\left(z_{t}|x_{1:t-1},z_{1:t-1},x_{1:L}\right)\Big{]}\Bigg{]}\Bigg{]}$
$\quad-\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}\left(z_{t+1:L+T}|x_{1:L+T}\right)}\Bigg{[}\mathbb{E}_{q_{\varphi}\left(z_{1:t-1}|x_{1:L+T}\right)}\Bigg{[}\mathbb{E}_{q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)}\Big{[}\log p_{\theta}\left(x_{t}|x_{1:t-1},z_{1:t},x_{1:L}\right)\Big{]}\Bigg{]}\Bigg{]}+\log p\left(x_{1:L+T}|x_{1:L}\right)$
$=\,$ $\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}\left(z_{t+1:L+T}|x_{1:L+T}\right)}\Bigg{[}\mathbb{E}_{q_{\varphi}\left(z_{1:t-1}|x_{1:L+T}\right)}\Big{[}\mathrm{KL}\left(q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)\,\middle\|\,p_{\theta}\left(z_{t}|x_{1:t-1},z_{1:t-1},x_{1:L}\right)\right)\Big{]}\Bigg{]}$
$\quad-\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}\left(z_{t+1:L+T}|x_{1:L+T}\right)}\Bigg{[}\mathbb{E}_{q_{\varphi}\left(z_{1:t-1}|x_{1:L+T}\right)}\Bigg{[}\mathbb{E}_{q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)}\Big{[}\log p_{\theta}\left(x_{t}|x_{1:t-1},z_{1:t},x_{1:L}\right)\Big{]}\Bigg{]}\Bigg{]}+\log p\left(x_{1:L+T}|x_{1:L}\right)$
$=\,$ $\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}\left(z_{t+1:L+T}|x_{1:L+T}\right)}\Bigg{[}\mathbb{E}_{q_{\varphi}\left(z_{1:t-1}|x_{1:L+T}\right)}\Bigg{[}\mathrm{KL}\left(q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)\,\middle\|\,p_{\theta}\left(z_{t}|x_{1:t-1},z_{1:t-1},x_{1:L}\right)\right)$
$\qquad\quad-\mathbb{E}_{q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)}\left[\log p_{\theta}\left(x_{t}|x_{1:t-1},z_{1:t},x_{1:L}\right)\right]\Bigg{]}\Bigg{]}+\log p\left(x_{1:L+T}|x_{1:L}\right)$
Rearranging the above equation:
$\log p\left(x_{1:L+T}|x_{1:L}\right)=\,$ ${\rm KL}\left(q_{\varphi}\left(z_{1:L+T}|x_{1:L+T}\right)\,\middle\|\,p\left(z_{1:L+T}|x_{1:L+T}\right)\right)$
$+\sum_{t=1}^{L+T}\mathbb{E}_{q_{\varphi}\left(z_{t+1:L+T}|x_{1:L+T}\right)}\Biggl{[}\mathbb{E}_{q_{\varphi}\left(z_{1:t-1}|x_{1:L+T}\right)}\Biggl{[}\mathbb{E}_{q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)}\Biggl{[}\log p_{\theta}\left(x_{t}|x_{1:t-1},z_{1:t},x_{1:L}\right)\Biggr{]}$
$\qquad\qquad\quad-\mathrm{KL}\left(q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)\,\middle\|\,p_{\theta}\left(z_{t}|x_{1:t-1},z_{1:t-1},x_{1:L}\right)\right)\Biggr{]}\Biggr{]}$
Since ${\rm{KL}}\left({{q_{\varphi}}\left({{z_{1:L+T}}|{x_{1:L+T}}}\right)||p\left({{z_{1:L+T}}|{x_{1:L+T}}}\right)}\right)$ is non-negative, we can derive the ELBO as follows:
$\log p\left(x_{1:L+T}|x_{1:L}\right)\geq L\left(\theta,\phi;x_{1:L+T}\right)=\sum_{t=1}^{L+T}\Bigg{(}$ $\mathbb{E}_{q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)}\left[\log p_{\theta}\left(x_{t}|x_{1:t-1},z_{1:t},x_{1:L}\right)\right]$
$-\mathrm{KL}\left(q_{\varphi}\left(z_{t}|z_{1:t-1},x_{1:L+T}\right)\|p_{\theta}\left(z_{t}|x_{1:t-1},z_{1:t-1},x_{1:L}\right)\right)\Bigg{)}$
Finally, based on the assumptions in Eq. (1) and (2), the ELBO of StoxLSTM can be derived as:
$\log p\left(x_{1:L+T}|x_{1:L}\right)$ $\geq L\left(\theta,\phi;x_{1:L+T}\right)=$
$\sum_{t=1}^{L}\Bigg{(}\mathbb{E}_{q_{\varphi}\left(z_{t}|z_{t-1},x_{1:L+T}\right)}\left[\log p_{\theta}\left(x_{t}|z_{t},x_{1:t-1}\right)\right]-\mathrm{KL}\left(q_{\varphi}\left(z_{t}|z_{t-1},x_{1:L+T}\right)\|p_{\theta}\left(z_{t}|z_{t-1},x_{1:t-1}\right)\right)\Bigg{)}$
$+\sum_{t=L+1}^{L+T}\Bigg{(}\mathbb{E}_{q_{\varphi}\left(z_{t}|z_{t-1},x_{1:L+T}\right)}\left[\log p_{\theta}\left(x_{t}|z_{t},x_{1:L}\right)\right]-\mathrm{KL}\left(q_{\varphi}\left(z_{t}|z_{t-1},x_{1:L+T}\right)\|p_{\theta}\left(z_{t}|z_{t-1},x_{1:L}\right)\right)\Bigg{)}$
Appendix B Supplementary Experiments
B.1 CRPS Evaluation Metric
Generative time series forecasting can also use the Continuous Ranked Probability Score (CRPS) to evaluate the prediction results. CRPS is computed by comparing the ground truth with the cumulative distribution function (CDF):
[…]
For point forecasting models, the point prediction can be treated as a degenerate CDF. Substituting into the formula and simplifying yields:
[…]
indicating that CRPS for point forecasts is equivalent to MAE.
Table 5: Long-term time series forecasting results. The best results are in bold and the second best are in underline. The evaluation metric is CRPS.
Models StoxLSTM xLSTM-Mixer xLSTMTime P-sLSTM iTransformer PatchTST FEDformer Informer Autoformer Dlinear TimeMixer++
Weather
96 0.167 0.184 0.187 0.208 0.214 0.198 0.296 0.384 0.336 0.237 0.205
192 0.206 0.226 0.236 0.256 0.254 0.241 0.336 0.544 0.367 0.282 0.245
336 0.225 0.266 0.272 0.297 0.296 0.282 0.380 0.523 0.395 0.319 0.265
720 0.254 0.323 0.326 0.350 0.347 0.334 0.428 0.741 0.428 0.362 0.334
Electricity
96 0.223 0.218 0.221 0.226 0.240 0.222 0.308 0.368 0.317 0.237 0.222
192 0.242 0.235 0.243 0.243 0.253 0.240 0.315 0.386 0.334 0.249 0.235
336 0.253 0.250 0.259 0.262 0.269 0.259 0.329 0.394 0.338 0.267 0.245
720 0.270 0.276 0.276 0.293 0.317 0.290 0.355 0.439 0.331 0.301 0.310
Solar
96 0.178 0.313 0.306 0.232 0.237 0.224 0.383 0.247 0.446 0.377 0.231
192 0.212 0.328 0.331 0.241 0.261 0.251 0.364 0.251 0.561 0.397 0.263
336 0.223 0.330 0.336 0.248 0.273 0.252 0.397 0.287 0.588 0.415 0.269
720 0.241 0.324 0.324 0.249 0.275 0.269 0.402 0.301 0.633 0.412 0.270
ETTh1
96 0.396 0.386 0.395 0.405 0.405 0.400 0.419 0.713 0.459 0.399 0.403
192 0.407 0.417 0.416 0.431 0.436 0.429 0.448 0.792 0.482 0.416 0.441
336 0.412 0.429 0.437 0.458 0.458 0.440 0.465 0.809 0.496 0.443 0.434
720 0.421 0.448 0.465 0.512 0.491 0.468 0.507 0.865 0.512 0.490 0.451
ETTh2
96 0.305 0.329 0.333 0.378 0.349 0.337 0.388 1.525 0.397 0.353 0.328
192 0.337 0.375 0.378 0.442 0.400 0.382 0.439 1.931 0.452 0.418 0.379
336 0.354 0.401 0.403 0.494 0.432 0.384 0.487 1.835 0.486 0.465 0.398
720 0.382 0.424 0.430 0.669 0.445 0.422 0.474 1.625 0.511 0.551 0.415
ETTm1
96 0.307 0.328 0.335 0.343 0.368 0.346 0.419 0.571 0.475 0.343 0.334
192 0.336 0.354 0.361 0.369 0.391 0.370 0.441 0.669 0.496 0.365 0.362
336 0.356 0.374 0.379 0.391 0.420 0.392 0.459 0.871 0.537 0.386 0.391
720 0.384 0.407 0.411 0.424 0.459 0.420 0.490 0.823 0.561 0.421 0.423
ETTm2
96 0.222 0.244 0.250 0.302 0.264 0.256 0.287 0.453 0.339 0.260 0.245
192 0.256 0.285 0.288 0.376 0.309 0.296 0.328 0.563 0.340 0.303 0.291
336 0.283 0.322 0.322 0.438 0.348 0.329 0.366 0.887 0.372 0.342 0.343
720 0.347 0.377 0.380 0.482 0.407 0.385 0.415 1.388 0.419 0.421 0.399
Traffic
96 0.257 0.236 0.242 0.332 0.268 0.249 0.366 0.391 0.388 0.282 0.253
192 0.310 0.241 0.253 0.342 0.276 0.256 0.373 0.379 0.382 0.287 0.258
336 0.320 0.250 0.261 0.345 0.283 0.264 0.383 0.420 0.337 0.296 0.263
720 0.335 0.283 0.287 0.357 0.302 0.286 0.382 0.472 0.408 0.315 0.282
ILI
24 0.579 0.707 0.694 1.309 0.883 0.754 1.260 1.449 1.182 1.081 0.823
36 0.680 0.712 0.722 1.330 0.870 0.870 1.080 1.463 1.038 0.963 0.835
48 0.710 0.721 0.725 1.326 0.815 0.815 1.078 1.542 1.145 1.024 0.818
60 0.768 0.737 0.715 1.384 0.788 0.788 1.157 1.543 1.114 1.096 0.839
B.2 Additional Experimental Results
We evaluate StoxLSTM’s CRPS across multiple datasets and prediction horizons in long-term forecasting tasks. For point prediction models, MAE and CRPS results coincide. We also assess CRPS for all baseline models. The results in Table 5 demonstrate that StoxLSTM achieves the lowest CRPS on most datasets, with xLSTM-Mixer ranking second. For instance, on the Solar dataset, StoxLSTM attains an average CRPS of 0.2135, which is 34.0% lower than the 0.3238 average CRPS of xLSTM-Mixer. Similarly, on the ETTm2 dataset, StoxLSTM achieves an average CRPS of 0.277, outperforming xLSTM-Mixer’s 0.307 by 9.8%. These results indicate that StoxLSTM provides more accurate probabilistic forecasts across diverse datasets.